# Enhancing Model Context Protocol (MCP) with Context-Aware Server Collaboration

### v1

@@ 4 Empirical Evaluation
To validate the advantages of our proposed CA-MCP, we evaluate its performance against the traditional MCP baseline on two benchmarks: (i) the TravelPlanner [14] benchmark for real-world itinerary planning, and (ii) the REALM-Bench [15] Wedding Logistics (P5) scenario for multi-agent scheduling. We report metrics that capture execution efficiency, constraint satisfaction, semantic accuracy, and coordination effectiveness.
@@ 4.1 Use Case 1: Travel Planning (TravelPlanner)
The TravelPlanner [14] benchmark provides over 1,200 multi-turn tasks that require generating realistic itineraries under budgetary, temporal, and preference constraints. We sample 500 queries such as: “Plan a three-day trip around Seattle with adventurous activities, vegan options, and a \$1500 budget.” Each task requires integrating location recommendations, weather forecasts, hotel bookings, and dining suggestions.
@@
Figure 3: Execution Time (sec) on TravelPlanner [14] (500 queries). (a) Overlaid histograms for Traditional (blue) and Context-Aware (green); dashed lines show means: Traditional = 41.989, Context-Aware = 13.523. (b) Pairwise differences (Traditional – Context-Aware); positive values indicate Context-Aware is faster. Paired t-test: mean_diff = 28.465s, $\mathtt{sd}=18.059s$, $n=500$, $\mathtt{pval}=1.60e{-137}$. Takeaway: CA-MCP markedly reduces latency.
Figure 4: Completeness Score on TravelPlanner [14] (500 queries). (a) Overlaid histograms for Traditional (blue) and Context-Aware (green); dashed lines show means: Traditional = 0.764, Context-Aware = 1.000. (b) Pairwise differences (Context-Aware – Traditional); positive values indicate higher completeness for Context-Aware. Paired t-test: mean_diff = 0.236, $\mathtt{sd}=0.2408$, $n=500$, $\mathtt{pval}=4.39e{-75}$. Takeaway: CA-MCP consistently attains higher completeness.
Figure 5: BERTScore (F1) on TravelPlanner [14] (500 queries). (a) Overlaid histograms for Traditional (blue) and Context-Aware (green); dashed lines show means: Traditional = 0.745, Context-Aware = 0.757. (b) Pairwise differences (Context-Aware – Traditional). Paired t-test: mean_diff = 0.012, $\mathtt{sd}=0.04728$, $n=500$, $\mathtt{pval}=2.35e{-08}$. Takeaway: CA-MCP yields a modest but consistent lift in semantic accuracy.
Figure 6: ROUGE-L on TravelPlanner [14] (500 queries). (a) Overlaid histograms for Traditional (blue) and Context-Aware (green); dashed lines show means: Traditional = 0.031, Context-Aware = 0.027. (b) Pairwise differences (Context-Aware – Traditional). Paired t-test: mean_diff = 0.004-0.004, $\mathtt{sd}=0.01173$, $n=500$, $\mathtt{pval}=1.24e{-13}$. Takeaway: Differences are small, with a slight edge for Traditional.
@@ Results.
Metric Traditional MCP CA-MCP Improvement
@@
Table 2: TravelPlanner [14] results comparing Traditional and CA-MCP.
Thus, Figure 3 shows that a CA-MCP can provide significant speed advantages relative to a traditional MCP. In our experiments, we deliberately use the most conceptually straightforward realizations of the Traditional MCP (Figure 1) and the CA-MCP (Figure 2). It is likely that alternative (e.g., multithreaded or dynamic) implementations of a Traditional MCP could outperform our simple Context-Aware realization; likewise, software and hardware optimizations could further accelerate the CA-MCP. We leave such engineering explorations to future work. The goal of the present study is to demonstrate, scientifically, that the conceptual idea of a CA-MCP improves efficiency over a straightforward conceptual implementation of a Traditional MCP.
@@ 4.2 Use Case 2: Wedding Logistics (REALM-Bench P5)
REALM-Bench [15] is a rigorous benchmark for evaluating LLM-based multi-agent planning and scheduling in dynamic, real-world scenarios. We adapt problem P5 (Wedding Reunion), which requires orchestrating guest arrivals, errands, and shared vehicles under deadline and capacity constraints. The scenario challenges stateless LLM orchestration due to interdependencies and timing conflicts.
@@ Results.
Metric Traditional MCP CA-MCP Improvement
@@
Coordination Score (0/1) 0 1 Improved coordination
@@
Table 3: REALM-Bench Wedding Logistics (P5) [15] results comparing Traditional and CA-MCP. CA-MCP achieves superior efficiency and coordination with fewer central LLM calls
@@ 5 Discussion
  • LLM Call Reduction: CA-MCP consistently required fewer central LLM invocations, mitigating context-window bottlenecks [13].
@@ 6 Future Work
  • Parallel and Asynchronous Execution: Enhancing the SCS to support true asynchronous server execution, enabling large-scale parallel planning.
  • Server-Level Learning: Allowing servers to adapt their interaction strategies based on historical execution traces, further reducing dependence on the central LLM.
  • Multimodal Integration: Extending the framework to include non-textual servers (e.g., vision, audio, structured data) contributing to the shared context.
@@ 7 Conclusion
This paper introduced the CA-MCP, an extension of the Model Context Protocol that reallocates the roles and responsibilities between a central LLM and specialized MCP servers. Our use of a shared context via the SCS could enable autonomous server coordination, reduced redundant LLM calls, and improved execution efficiency and robustness across benchmarks.
Empirical evaluation on TravelPlanner and REALM-Bench suggests that CA-MCP could deliver substantial improvements in latency (up to 73.5% faster), constraint satisfaction (perfect completeness scores), and coordination efficiency (45.5% shorter makespan). These results suggest that shifting orchestration from the central LLM to shared state can be promising, though broader assessment across tasks is needed to fully characterize performance and generality.
Together, these results establish CA-MCP as a practical and scalable design for LLM-driven multi-agent systems, bridging the gap between centralized orchestration and distributed autonomy.

### v2

@@ 4 Empirical Evaluation
To validate the advantages of our proposed CA-MCP, we evaluate its performance against the traditional MCP baseline on two benchmarks: (i) the TravelPlanner (Yang et al., 2024) benchmark for real-world itinerary planning, and (ii) the REALM-Bench (Geng and Chang, 2025) Wedding Logistics (P5) scenario for multi-agent scheduling. We report metrics that capture execution efficiency, constraint satisfaction, semantic accuracy, and coordination effectiveness.
@@ 4.1 Use Case 1: Travel Planning (TravelPlanner)
The TravelPlanner (Yang et al., 2024) benchmark provides over 1,200 multi-turn tasks that require generating realistic itineraries under budgetary, temporal, and preference constraints. We sample 500 queries such as: “Plan a three-day trip around Seattle with adventurous activities, vegan options, and a \$1500 budget.” Each task requires integrating location recommendations, weather forecasts, hotel bookings, and dining suggestions.
Evaluation Metrics.
We measure: Execution Time (wall-clock seconds), Completeness (0–1, fraction of pipeline stages completed), BERTScore F1 (Zhang et al., 2020) (semantic similarity), ROUGE-L (Lin, 2004) (sequence overlap), and LLM Calls (orchestration overhead).
@@
Figure 2: Execution Time (sec) on TravelPlanner (Yang et al., 2024) (500 queries). (a) Overlaid histograms for Traditional (blue) and Context-Aware (green); dashed lines show means: Traditional = 41.989, Context-Aware = 13.523. (b) Pairwise differences (Traditional – Context-Aware); positive values indicate Context-Aware is faster. Paired t-test: mean_diff = 28.465s, $\mathtt{sd}=18.059s$, $n=500$, $\mathtt{pval}=1.60e{-137}$. Takeaway: CA-MCP markedly reduces latency.
Figure 3: Completeness Score on TravelPlanner (Yang et al., 2024) (500 queries). (a) Overlaid histograms for Traditional (blue) and Context-Aware (green); dashed lines show means: Traditional = 0.764, Context-Aware = 1.000. (b) Pairwise differences (Context-Aware – Traditional); positive values indicate higher completeness for Context-Aware. Paired t-test: mean_diff = 0.236, $\mathtt{sd}=0.2408$, $n=500$, $\mathtt{pval}=4.39e{-75}$. Takeaway: CA-MCP consistently attains higher completeness.
Figure 4: BERTScore (F1) on TravelPlanner (Yang et al., 2024) (500 queries). (a) Overlaid histograms for Traditional (blue) and Context-Aware (green); dashed lines show means: Traditional = 0.745, Context-Aware = 0.757. (b) Pairwise differences (Context-Aware – Traditional). Paired t-test: mean_diff = 0.012, $\mathtt{sd}=0.04728$, $n=500$, $\mathtt{pval}=2.35e{-08}$. Takeaway: CA-MCP yields a modest but consistent lift in semantic accuracy.
Figure 5: ROUGE-L Score on TravelPlanner (Yang et al., 2024) (500 queries). (a) Overlaid histograms for Traditional (blue) and Context-Aware (green); dashed lines show means: Traditional = 0.031, Context-Aware = 0.027. (b) Pairwise differences (Context-Aware – Traditional). Paired t-test: mean_diff = 0.004-0.004, $\mathtt{sd}=0.01173$, $n=500$, $\mathtt{pval}=1.24e{-13}$. Takeaway: Differences are small, with a slight edge for Traditional.
@@ Results.
Table 2: TravelPlanner (Yang et al., 2024) results comparing Traditional and CA-MCP.
Metric Traditional CA-MCP Improvement
MCP
@@
Thus, Figure 2 shows that a CA-MCP can provide significant speed advantages relative to a traditional MCP.
@@ 4.2 Use Case 2: Wedding Logistics (REALM-Bench P5)
REALM-Bench (Geng and Chang, 2025) is a rigorous benchmark for evaluating LLM-based multi-agent planning and scheduling in dynamic, real-world scenarios. We adapt problem P5 (Wedding Reunion), which requires orchestrating guest arrivals, errands, and shared vehicles under deadline and capacity constraints. The scenario challenges stateless LLM orchestration due to interdependencies and timing conflicts.
Evaluation Metrics.
Following REALM-Bench (Geng and Chang, 2025), we measure: Goal Satisfaction (0–1, objectives achieved), Constraint Satisfaction (0–1, hard constraints met), Makespan (minutes, schedule duration), Coordination Score (0/1, coordination effectiveness or batching indicator), LLM Calls, and Execution Time (seconds).
@@ Results.
Table 3: REALM-Bench Wedding Logistics (P5) (Geng and Chang, 2025) results comparing Traditional and CA-MCP. CA-MCP achieves superior efficiency and coordination with fewer central LLM calls
Metric Traditional CA-MCP Improvement
MCP
@@
Coordination Score (0/1) 0 1 Improved
@@ 5 Discussion
  • LLM Call Reduction: CA-MCP consistently required fewer central LLM invocations, mitigating context-window bottlenecks (Agashe et al., 2025).
@@ 6 Future Work and Limitations
  • Parallel and Asynchronous Execution: Enhancing the SCS to support asynchronous server execution, enabling large-scale parallel planning.
  • Server-Level Learning: Allowing servers to adapt interaction strategies based on historical execution traces, further reducing dependence on the central LLM.
  • Multimodal Integration: Extending the framework to include non-textual servers (e.g., vision, audio, structured data) contributing to shared context.
@@
Limitations.
In our experiments (Tables 13 and Figures 25), we used straightforward realizations of Traditional MCP (Figure 1(a)) and CA-MCP (Figure 1(b)) to isolate the specific benefit of the SCS without hardware or software optimizations such as parallel multithreading. These optimizations could improve performance of both approaches and change their relative ordering. We leave such engineering explorations to future work. The goal of this study is to demonstrate that the conceptual idea of CA-MCP improves efficiency over a straightforward implementation of Traditional MCP.
@@ 7 Conclusion
This paper introduced CA-MCP, an extension of the Model Context Protocol that reallocates roles between a central LLM and specialized MCP servers. Our SCS enables autonomous server coordination, reducing redundant LLM calls and improving execution efficiency.
Empirical evaluation on TravelPlanner and REALM-Bench suggests that CA-MCP shows substantial improvements: up to 73.5% faster execution, perfect completeness scores, and 45.5% shorter makespan. These results suggest that shifting orchestration from the central LLM to shared state can be promising, though broader assessment is needed to fully characterize performance and generality.
Together, these results establish CA-MCP as a practical and scalable design for LLM-driven multi-agent systems, bridging centralized orchestration and distributed autonomy.
Impact Statement
This paper advances multi-agent LLM orchestration through architectural improvements reducing computational overhead by 67.8% (execution time) and 60% (LLM calls). These gains could reduce deployment costs and environmental impact of AI systems. As with any AI advancement, improved efficiency could enable both beneficial and harmful applications; deployment should include appropriate safeguards and domain-specific ethical review. This foundational research does not introduce novel ethical concerns beyond those inherent to advancing AI agent capabilities.